OpenStudy (anonymous):

\[x^{4}-30x ^{2}+125=0\]

OpenStudy (zzr0ck3r):

\(x^4-25x^2-5x^2+125=x^2(x^2-25)-5(x^2-25)=(x^2-5)(x^2-25)\)

OpenStudy (zzr0ck3r):

can you go any further?

OpenStudy (anonymous):

i wanna make sure i have it right

OpenStudy (anonymous):

my teacher put make an appropiate substition

OpenStudy (zzr0ck3r):

Let \(u=x^2\) then you have \(u^2-30u+125=0\iff (u-5)(u-25)=0\iff(x^2-5)(x^2-25)=0\)